(-3.2*10 to the power of -7)(-8.6*10 to the power of -8)
Answer by CubeyThePenguin(3113)

(-3.2 * 10^(-7)) * (-8.6 * 10^(-8)) = 27.52 * 10^(-15)
= 2.752 * 10^(-14)